OpenStudy (lifeisadangerousgame):

f(x) = (x+ 1)^3/4 For the function f, defined above, what are all the values of x for which f(x) is a real number? The answer is x >= -1, why?

OpenStudy (irishboy123):

\(\large f(x) = (x+ 1)^{\frac{3}{4}} = ((x+ 1)^{3} )^{\frac{1}{2}})^\frac{1}{2}\) so put in x = -1.01 and work out f(x)

OpenStudy (lifeisadangerousgame):

Ahh, I see! Ah shoot, that was an easy one .-.

OpenStudy (lifeisadangerousgame):

Thank you ^_^